### v1

# ETH Zurich

Abstract
Annotation and classification of legal text are central components of empirical legal research. Traditionally, these tasks are often delegated to trained research assistants. Motivated by the advances in language modeling, empirical legal scholars are increasingly turning to prompting commercial models, hoping that it will alleviate the significant cost of human annotation. Despite growing use, our understanding of how to best utilize large language models for legal tasks remains limited. We conduct a comprehensive study of 260 legal text classification tasks, nearly all new to the machine learning community. Starting from GPT-4 as a baseline, we show that it has non-trivial but highly varied zero-shot accuracy, often exhibiting performance that may be insufficient for legal work. We then demonstrate that a lightly fine-tuned Llama 3 model vastly outperforms GPT-4 on almost all tasks, typically by double-digit percentage points. We find that larger models respond better to fine-tuning than smaller models. A few tens to hundreds of examples suffice to achieve high classification accuracy. Notably, we can fine-tune a single model on all 260 tasks simultaneously at a small loss in accuracy relative to having a separate model for each task. Our work points to a viable alternative to the predominant practice of prompting commercial models. For concrete legal tasks with some available labeled data, researchers are better off using a fine-tuned open-source model.
1 Introduction
The legal system generates a staggering volume of complex documents. United States federal courts alone process hundreds of thousands of cases a year, each having substantial case files. Much empirical legal research involves the systematic collection and analysis of such data in order to understand how laws function in practice and what impact they have on society. What limits researchers across the board is the cost of annotating and classifying legal documents. Legal classification tasks vary in complexity, but often require substantial expertise and effort. Employing trained research assistants stretches to a few hundred, perhaps a few thousand documents at a time, but is no match for the sheer scale of legal data.
Motivated by the rapid advances in large language models, law scholars increasingly try out commercial models, such as GPT-4, on a variety of legal tasks, hoping to boost the efficiency of legal research. For lack of an alternative, researchers access these models in a zero-shot manner [^1] Fine-tuning GPT-4 is currently unavailable to all API users and requires special access.. The underlying assumption is that prompting GPT-4 provides the best solution to the problem that is currently available. In this work, we critically examine this assumption.
1.1 Our contributions
We introduce and study a collection of 260 legal classification tasks, nearly all new to the machine learning community. The tasks we introduce are actual legal classification tasks based on the U.S. Supreme Court (Spaeth et al., 2023) and Court of Appeals (Songer, ) databases that provide rich annotations for court files. We use these annotations as labels to derive challenging multi-class machine learning tasks. In particular, our suite of tasks significantly extends and strengthens valuable recent efforts to benchmark language models for legal tasks (Guha et al., 2023).
Our primary finding is that fine-tuning a single Llama 3 8B Inst (MetaAI, 2024) model on all classification tasks achieves vastly superior performance to GPT-4222We evaluate gpt-4-0613, which is what at the time of writting gpt-4 points to in the OpenAI API. The recently released GPT-4o model is currently not available for our region via the Azure OpenAI Service. zero-shot (Figure 1). Although it is expected that fine-tuning helps, the strong superiority of fine-tuning an open source model at much smaller scale is surprising. After all, GPT-4 is orders of magnitude larger according to available information and interpretation. Based on our comprehensive evaluation, we argue that the zero-shot performance of GPT-4 is far from sufficient for actual legal work. In contrast, a fine-tuned model provides a simple and practical alternative with much higher accuracy. In more detail, our primary contributions the following:
Figure 1: The cost of zero-shot: Difference in accuracy between Lawma 8B and GPT4. Each vertical bar represents the accuracy difference on one task, sorted in ascending order.
  • To establish baselines, we perform a zero-shot evaluation of recent state-of-the-art open source models (Figure 3), including Llama 3 8B Inst, Llama 3 70B Inst, Mistral 7B Inst, Saul 7B Inst, and Mixtral 8x7B Inst, and the commercial model GPT-4. Among these, only Llama 3 70B Inst and GPT-4 achieve zero-shot performance significantly better than a constant classifier that outputs the majority class. Still, there are dozens of tasks were both models perform worse than random guessing (Figure 4). Averaged across all tasks, GPT-4 has zero-shot accuracy 62.9% compared with 58.4% for Llama 3 70B. The constant classifier that outputs the majority class for each task achieves 41.7% accuracy, a number determined by class imbalance and the number of classes in each task. GPT-4 and Llama 3 achieve similar performance on the Court of Appeals tasks (63.4% and 60.3%, respectively), whereas GPT-4 has significantly higher accuracy on the Supreme Court tasks (59.8% versus 47.1%). Few-shot prompting GPT-4 does not improve performance. Section 2.3 provides more details.
  • Fine-tuning a single Llama 3 8B Instruct model, which we call Lawma 8B, outperforms GPT-4 zero-shot an all but a few of tasks. The improvements are typically in the double digit accuracy points, specifically, 22.6 percentage points on average for Supreme Court tasks and 16.5 points on average over Appeals Court tasks. Figure 1 displays the accuracy deficit of zero-shot prompting relative to fine-tuning across all tasks. In absolute terms, Lawma 8B achieves 82.4% accuracy on Supreme Court tasks and 79.9% accuracy on Appeals Court tasks.
    We additionally fine-tune a single Llama 3 70B Instruct model, which we call Lawma 70B, and obtain further small single digit improvements in performance. Lawma 70B outperforms Lawma 8B on average by 1.7 percentage points for Supreme Court tasks and 1.6 points for Appeals Court tasks.
  • We systematically evaluate the effect of finetuning on model performance at different scales. Larger models respond better to fine-tuning than smaller models. Across nine different base models, mean task accuracy increases steadily with pretraining compute as Figure 7 illustrates. However, we observe signs of diminishing returns. For Appeals Court tasks, scaling fine-tuning from Pythia 1B to Llama 3 70B yields an average improvement of only 8.5 accuracy points, despite a 3000x increase in pretraining compute. This suggests that, in the future, major improvements may not come from model scale alone.
  • Fine-tuning is data efficient. A few hundred examples typically suffice to achieve high classification accuracy. This is crucial, since labeling a few hundred data points is often financially feasible for many legal scholars (Hall and Wright, 2008), whereas labeling many thousands may not. Figure 8 shows how training accuracy increases with the number of training examples.
  • We can simultaneously fine-tune on all 260 tasks. There is not a large loss compared with fine-tuning on a specific task (Figure 9). This is desirable in practice, as it obviates the need to train and maintain a separate model for each task.
  • Fine-tuning generalizes to unseen tasks. We show that fine-tuning Llama 3 8B Inst only on the Court of Appeals tasks improves its average accuracy on Supreme Court tasks by 18.8 accuracy points (Figure 10).
  • We contextualize our accuracy numbers with intercoder agreement rates. Our analysis reveals task heterogeneity in the relationship between model accuracy and intercoder agreement.
Our results speak to the power of specialization for legal classification tasks. They challenge the prevailing narrative about the suitability of the zero-shot paradigm. Fine-tuning provides a simple, practical, and far more accurate solution than zero-shot prompting. These insights suggest that the empirical legal community should invest in an ecosystem of fine-tuned models for relevant classification tasks. From a benchmarking perspective, our tasks provide non-trivial measurements of model performance that may be of independent interest.
1.2 Related work and background
Adoption of large language models in the legal community.
The legal community has moved relatively quickly in adopting GPT models. Several startups have begun using incorporating large language models, including GPT, into legal products (Wiggers, 2022). Lexis Nexis, a major commercial provider of law-related services, has partnered with Open AI and Anthropic to offer legal text generation (LexisNexis, 2023). Legal scholars have evaluated GPT’s performance on the bar exam (Katz et al., 2024) as well as law school exam (Choi et al., 2023). Choi and Schwarcz (2023) examined how GPT-4 can improve student performance on law school exams. Nay et al. (2024) examined how LLMs perform on answering multiple choice questions related to tax law. Gray et al. (2024) used GPT models to extract information from cases concerning the factors that predict the constitutionality of police stops. Choi (2023) used GPT-4 to extract information concerning interpretative techniques from U.S. Supreme Court decisions. Livermore et al. (2023) tested the performance of GPT models for categorizing cases by issue areas and in recommending citations based on case similarity. Savelka and Ashley (2023) evaluate the zero-shot performance of GPT-4 on a variety of semantic legal annotation tasks. Engel and Mcadams (2024) ask GPT for the ordinary meaning of statutory terms. In the area of corporate law, Frankenreiter and Talley (2024) use GPT-4 to extract information about the contents of corporate charters.
Benchmarks and models for legal tasks.
LegalBench (Guha et al., 2023) is a recent multi-task benchmark for natural language understanding in legal domains. As of writing, LegalBench consists of 162 tasks gathered from 40 contributors. LegalBench draws on numerous earlier benchmarking efforts in different legal domains, specifically, inference on contracts (Koreeda and Manning, 2021; Hendrycks et al., 2021), merger agreement understanding (Wang et al., 2023), identifying the legal holding of a case (Zheng et al., 2021), statutory reasoning (Holzenberger and Van Durme, 2021), privacy compliance and policy (Wilson et al., 2016; Zimmeck et al., 2019; Ravichander et al., 2019), and identifying unfair clauses in terms of service (Lippi et al., 2019). We extend and strengthen these valuable efforts to benchmark large language models in legal settings. The 260 tasks we study are complementary to those in LegalBench. We focus on core legal classification tasks based on the U.S. Supreme Court Database (Spaeth et al., 2023) and the U.S. Courts of Appeals database (Songer, ), relevant to the field empirical legal studies. These databases provide rich information about judicial decisions, panel compositions, judge attributions, and case characteristics, that we use to derive challenging classification tasks. Section 2 continues with this discussion in more detail. We did not evaluate our model on LegalBench, since our model is specialized to the Supreme Court and Appeals Court data.
Legal-BERT (Chalkidis et al., 2020) is a BERT-like transformer model that was pretrained on a few hundred thousand legal documents, including more than 100,000 documents from EU legislation, and more than 150,000 documents from the Case Law Access Project portal. SaulLM (Colombo et al., 2024) is a more recent model trained on a legal corpus of 30 billion tokens. SaulLM is based on the Mistral 7B model (Jiang et al., 2023) that incorporates additional pretraining and instruction fine-tuning specific to legal text. Our work is rather orthogonal, evaluating the strength of specialized models rather than building new general-purpose models.
Data annotation and labeling.
Hall and Wright (2008) provide an overview of the use of human annotators in empirical legal studies. Student coders have been deployed to extract a wide variety of features from legal data. Although student researchers are much less expensive than private attorneys, the costs can quickly become prohibitive. Depending on the size of the document and the complexity of the task, research assistants can label roughly dozens of examples per hour. Projects involving the labeling of hundreds of documents are financially feasible for many legal scholars, but projects involving many thousands of documents are largely impractical. In an example of a larger annotation effort, Frankenreiter et al. (2021) employed human coders to annotate several thousands of corporate charters. Using ChatGPT for a similar task, Frankenreiter and Talley (2024) estimated that employing human coders would have been approximately ten times more costly than their approach.
Data annotation and labeling also play a major role in machine learning benchmarks and applications, see, e.g., Aroyo and Welty (2015); Gray and Suri (2019); Hardt and Recht (2022) for background. Dorner and Hardt (2024) give an extended discussion about label quality and annotator disagreement in the context of machine learning benchmarks.
1.3 Limitations
Fine-tuning increases accuracy to about $80\%$ in our evaluation suite compared with around $60\%$ for zero-shot prompting. While we are rather certain that $60\%$ accuracy is insufficient for consequential legal work, we emphasize that $80\%$ is still far from perfect. In addition, the variance in accuracy across tasks, while lower than for zero-shot prompting, remains high. Although our work meets the ethical and technical recommendations by Kapoor et al. (2024) for “developers of legal AI”, we maintain caution about the use of large language models for consequential legal tasks. To which extent these models are suitable for use in specific applications requires additional substantive investigation. We add that the legal documents we consider are exclusively from either the U.S. Supreme Court or appellate courts in the United States. We cannot speak to how these results may change for tasks in other legal domains within United States or legal systems in other countries.
2 Legal classification tasks
In this work, we focus on legal classification tasks. Legal classification tasks range in complexity, from extremely simple tasks that require little specialized knowledge, to highly sophisticated tasks that involve substantial expertise and judgment. Simple tasks would include identifying the parties to a case or the general issue area—for example, whether a case dealt with family law or commercial contracts. More sophisticated tasks involve specific legal knowledge, familiarity with legal principles or discourse, and the ability to engage in nuanced analogical or conceptual reasoning. For example, labeling the ideological valence of a decision requires the annotator to understand how specific legal issues map onto contemporary political debates, while labeling the standard of review applied by an appellate court requires detailed knowledge of these standards as well as the ability to parse procedural history. Many legal doctrines are quite complicated, involving multipart tests, nuanced exceptions, and balancing inquiries. Extracting features concerning such doctrines can lead to disagreement even among experienced annotators with considerable legal expertise.
More efficient ways to solve legal classification tasks would be tremendously useful in practice. A well functioning system to automatically extract relevant features from legal texts could, in particular, facilitate empirical legal study across a wide range of domains. This research could include not only social scientific study of the causes or consequences of judicial decisions, but also more traditional research modalities based on doctrinal interpretation (Livermore and Rockmore, 2019). There is an almost unlimited variety of features that legal scholars could study, ranging from the factors cited by judges when deciding the outcomes of property law disputes to the relationship between the party affiliation of judges and their use of different interpretative styles. With the digitization of legal texts at the U.S. state level and outside the United States, low-cost and flexible featurization can also boost efforts to show the geographic diffusion of legal concepts or approaches.
To summarize, our reasons to study legal classification tasks are both technical and substantive. From a technical machine learning perspective, these tasks provide highly non-trivial classification problems where even the best models leave much room for improvement. From a substantive legal perspective, efficient solutions to such classification problems have rich and important applications in legal research, as we discussed.
Figure 2: General statistics of the court opinions and legal classification tasks considered.
2.1 Data sources
Central to our study are the U.S. Supreme Court Database (Spaeth et al., 2023) (SCDB) and the U.S. Courts of Appeals database (Songer, ) (USCAD). The SCDB compiles comprehensive information on U.S. Supreme Court decisions from 1946 onward. Developed by Harold Spaeth, it includes variables such as case outcomes, issue areas, legal provisions, and vote counts. The USCAD contains detailed information about decisions made by the U.S. Courts of Appeals from 1925 to 1988. It includes data on judicial decisions, panel compositions, and case characteristics. Both databases provide essential tools for scholars conducting quantitative analyses of the judicial system, decision-making, ideological trends, and the impact of various factors on case outcomes.
The SCDB and USCAD have been instrumental in advancing research on judicial decision making within the fields of political science and empirical legal studies (Epstein et al., 2013; Segal and Spaeth, 2002; Martin and Quinn, 2002). These datasets have been used to drive a substantial research program by allowing scholars to systematically analyze large numbers of court cases, uncovering patterns, trends, and factors influencing judicial outcomes. By providing detailed information on case characteristics, judge attributes, and decision outcomes, these databases have enabled researchers to test theories of judicial behavior, examine the impact of ideology on court decisions, and explore the dynamics of judicial decision-making at different levels of the court system. The insights gained from research using these databases have had significant implications for legal practitioners, policymakers, and the broader legal community, contributing to a better understanding of how courts operate and how legal outcomes are shaped.
2.2 Construction of classification tasks
We construct a set of classification tasks using the variables of the Supreme Court and the U.S. Courts of Appeals databases. We construct a total of 260 distinct classification tasks, 38 of them corresponding to the Supreme Court database and 232 to the U.S. Court of Appeals. For each of these classification tasks, we construct a prompt template consisting of a general description of the task, followed by a multiple choice question containing each of the possible variable codes. We formulate the task description, question, and answer choices by closely following the language of the variable description of the databases’ documentation. Thereafter, for every case contained in these databases, we use the provided case citations to search for its corresponding majority opinion of the court on the Caselaw Access Project , a database of digitized court opinions.
We divide court opinions into a 70%/10%/20% train/validation/test split. Since many of the classification taks contain heavily imbalanced classes, we subsample the majority class such that there are at most as many task examples in the majority class as task examples in all other classes combined. Therefore, a constant classifier that outputs the majority class label will never achieve more than $50\%$ accuracy on any task. We find that on average over all tasks, the majority classifier has 41.7% accuracy. We plot some statistics of the tasks in Figure 2. First, court opinions tend to be long, with 12% having above 8,000 tokens, the typical maximum context size for current state-of-the-art models, such as Llama 3. Second, some tasks have a large number of classes, with 28% of tasks having more than 10 classes. Third, there is a large variability in terms of the number of task examples, ranging from a couple dozen to 18500 task examples. In the remainder of the article, we highlight ten different tasks, six from the SCDB and four from the USCAD. These tasks differ in complexity, ranging from relatively simple tasks (e.g., determining the issue area) to relatively complex ones (e.g., determining the ideological “direction” of the court decision). Section 2.4 contains a brief description of these tasks, Section B the full list of all tasks.
Example of the prompt template.
We use a prompt template identical to the one for the MMLU benchmark (Hendrycks et al., 2020). We provide as an example the prompt template corresponding to the Supreme Court precedent alteration variable (“sc_precedentalteration” in Section B).
What follows is an opinion from the Supreme Court of the United States.
Your task  is to identify  whether the opinion effectively says that
the decision in this case "overruled" one or more of the Court\’s own
precedents. Alteration also extends to language in the majority opinion
that states that a precedent of the Supreme Court has been "disapproved,"
or is "no longer good law". Note, however, that alteration does not
apply to cases in which the Court "distinguishes" a precedent.
[COURT OPINION]
 Question: Did the the decision of the court overrule one or more
 of the Court’s own precedents?
 A. Yes
 B. No
 Answer:
Due to diverse set of models and very large number of tasks under consideration, we decided to perform no prompt tuning. [^3] Multiple choice prompts are popular for language model evaluation because they only requires a single forward pass per example. Since many popular benchmarks are phrased as multiple-choice, recent instruction tuned models tend to do well for them. Note that more involved prompting strategies –e.g., chain-of-thought (Wei et al., 2022)– can yield better task performance but are substantially more expensive. For legal tasks specifically, the choice of prompt can have a significant effect in performance (Li et al., 2024). Prompt tuning requires task-specific domain knowledge and can be reasonably time consuming. Instead, we formulate the task description, question, and answer choices by closely following the language of the variable description of the databases’ documentation.
2.3 Zero-shot and few-shot evaluation baselines
Figure 3: Zero-shot accuracy of models on different legal classification tasks. Top: Mean accuracy over all cases. Middle: Mean task accuracy. Bottom: Accuracy in 10 representative tasks. Error bars indicate 90% Clopper-Pearson exact Binomial confidence intervals.
To establish baselines, we evaluate the zero-shot performance of different language models on all 260 classification tasks. Figure 3 presents the accuracies achieved in different task categories (all, Supreme Court, Court of Appeals), as well as ten representative tasks. We find that only Llama 3 70B Inst and GPT-4 are consistently better than the constant classifier that outputs the majority class. Figure 4 compares the zero-shot accuracy of GPT-4 and Llama 3 70B across all 260 tasks.
Figure 4: Difference in zero-shot accuracy between GPT4 and Llama 3 70B Instruct. Each vertical bar represents the accuracy difference on one task, sorted in ascending order.
Confidence intervals.
There are two natural ways to compute aggregate accuracies across multiple tasks. One is to compute the accuracy across all cases in the task collection. Tasks with fewer cases have lower weight under this average. The other is to compute the accuracy on each task separately and average those out. Since some tasks have far fewer documents than others (see Figure 2), the latter has higher variance. We indicate this by plotting the average error bars across tasks. Hence, the error bars for mean task accuracy are larger. The error bars for the case average are generally negligible, while several very small tasks have large error bars.
Table 1: Zero-shot and few-shot accuracies for different models
Model All tasks Supreme Court Court of Appeals
LegalBERT 24.58 13.59 26.45
Saul 7B Instruct 34.41 20.23 36.82
Mistral 7B Instruct 39.90 19.52 43.36
Majority class 41.72 31.48 43.45
Llama 3 8B Instruct 42.58 32.80 44.24
Mixtral 8x7B Instruct 43.22 24.37 46.42
Llama 3 70B Instruct 58.37 47.13 60.27
GPT-4 62.89 59.78 63.42
GPT-4 $32k$ $3$-shot 58.38 53.99 59.12
Multi-shot prompting.
We also evaluated whether multi-shot prompting would lead to improvements. As Figure 2 shows, many documents have thousands of tokens. The default context window for GPT-4 is 8,000 tokens. A version with 32,000 tokens is available at twice the cost. We evaluated the $32k$ version with $3$-shot prompting. This, however, did not yield any improvements over zero-shot prompting the standard model. See the last row in Table 1.
2.4 Highlighted tasks
Throughout this paper, as in Figure 3, we provide detailed results for ten tasks. Six of these tasks are from the SCDB, and four are from the USCAD. We selected tasks that we believe are particularly relevant to the legal community and chose tasks with varying levels of complexity, ranging from relatively simple (e.g., determining the issue area) to more complex (e.g., determining the ideological ’direction’ of the court decision).
Four tasks from the USCAD and all tasks from the SCDB were selected to form pairs, with each pair consisting of one task from the USCAD and one from the SCDB that capture similar concepts. It is important to note that, despite capturing broadly similar concepts, the precise formulation of the tasks might differ between the USCAD and the SCDB, making them less than perfectly comparable. In addition to the four pairs, we include two tasks from the SCDB that involve determining features of the decision reviewed by the Supreme Court on the basis of the Supreme Court opinion.
The following is a description of the task pairs:
  • SC Issue Area / Songer Gen Issue: These tasks capture the case’s issue area, requiring a determination of whether the case belongs to one of several broadly defined categories, such as criminal cases or First Amendment cases. These tasks are expected to be of relatively low complexity.
  • SC Case Source / Songer Case Origin: These tasks require identifying the court or adjudication body where the case was originally initiated before moving up the judicial hierarchy. Like the previous pair, these tasks are expected to be of relatively low complexity.
  • SC Disposition / Songer Treatment: These tasks involve determining how the deciding court treated the lower court opinion it reviewed, such as whether it affirmed or reversed the opinion. We consider these tasks to be of relatively low complexity.
  • SC Direction / Songer Direction: These tasks involve determining the ideological ’direction’ of the decision, specifically whether the decision supports a “conservative” or “liberal” outcome. We consider these tasks to be comparably complex.
  • SC LC Disposition / SC LC Direction: These tasks involve determining the disposition and ideological ’direction’ of the decision reviewed by the Supreme Court. As these tasks require analyzing features of another decision based on the text of the Supreme Court decision, we consider these tasks to be comparably complex.
3 Fine-tuning and the power of specialization
Figure 5: Zero-shot accuracy of four different models in comparison with the accuracy of (fine-tuned) Lawma 8B and 70B on different legal classification tasks. Top: Mean accuracy over all cases. Middle: Mean task accuracy. Bottom: Accuracy in 10 representative tasks. Error bars indicate 90% Clopper-Pearson exact Binomial confidence intervals.
In this section, we present a detailed analysis of fine-tuning on different tasks. We start by fine-tuning Llama 3 8B Inst and Llama 3 70B Inst on all tasks simultaneously, resulting in our Lawma 8B and Lawma 70B models. We then perform additional fine-tuning experiments highlighting different aspects, including the scaling behaviour of fine-tuning larger base models, the sample efficiency of fine-tuning, its generalization to unseen tasks and Courts, and the effect of single task specialization.
We first fine-tune Llama 3 8B Inst and Llama 3 70B Inst on all tasks simultaneously. We refer to these fine-tuned models as Lawma 8B and Lawma 70B, respectively. We fine-tune on the 260 classification tasks described in Section 2.2. The fine-tuning dataset contains a total of 1.96B tokens. For Lawma 3 8B, we fine-tune for 3 epochs. For Lawma 3 70B, we fine-tune for 1 epoch. We find that additional epochs hurt performance. See Appendix A for additional details regarding the model training.
We compare in Figure 5 the task accuracies of Lawma 8B and 70B to that of their respective base models Llama 3 8B Inst and Llama 3 70B Inst, as well as GPT-4. Fine-tuning leads to large improvements in average task accuracy: Lawma 8B outperforms Llama 3 8B Inst by 37.7 accuracy points and Lawma 70B outperforms Llama 3 70B by 23.5 accuracy points. Both Lawma 8B and Lawma 70B outperform GPT4, Lawma 8B on average by 17.3 accuracy points and Lawma 70B on average by 19.0 accuracy points.
Figure 6: Distribution of task performance across all tasks for Llama 3, GPT4, and Lawma.
Moreover, the Lawma models are the top performing models in each of the 10 highlighted tasks. In fact, the both Lawma 8B and Lawma 70B outperforms GPT4 in about 95% of all tasks, see Figure 1. Figure 6 further demonstrates the large effect of fine-tuning by showing the histogram of task accuracies achieved by Lawma 8B and Lawma 70B in comparison with Llama 3 8B Inst, Llama 3 70B Inst, and GPT4.
We find that Lawma 8B closely matches the performance of Lawma 70B. Specifically, Lawma 70B outperforms Lawma 8B in terms of mean case accuracy by only 0.9 accuracy points for Supreme Court tasks and by 1.2 accuracy points for Appeals Court tasks (Figure 5). In terms of mean task accuracy, the difference between Lawma 8B and Lawma 70B is not statistically significant. We find that by training the 8B model on more epochs than the 70B model, it attains a similar loss as the 70B model. Further training either model results in overfitting and thus performance degrades on average. This suggests that, for our fine-tuning dataset, further scaling model size (e.g., fine-tuning GPT4) is unlikely to yield major improvements. A more promising direction is to instead improve the diversity and quantity of the fine-tuning data. On the flip side, practitioners may choose to use Lawma 8B instead of the 70B model with little cost in model performance.
3.1 Performance after fine-tuning scales with pretraining compute
The performance of specialized models tends to scale with pretraining compute (Dominguez-Olmedo et al., 2024). We investigate how performance after fine-tuning scales with the pretraining compute of the base model. We fine-tune the following models for a single epoch: Pythia 70M, Pythia 160M, Pythia 410M, Pythia 1B, Pythia 2.8B, Pythia 6.9B (Biderman et al., 2023), Llama 2 7B (Touvron et al., 2023), Llama 3 8B Inst and Llama 3 70B Inst. We fine-tune on all 260 tasks simultaneously. We approximate pretraining compute in FLOPs as $C\approx 6\cdot N\cdot D$ (Kaplan et al., 2020), where $N$ is model size and $D$ is the number of pretraining tokens.
Figure 7: Performance after one epoch of fine-tuning increases monotonically in the amount of pretraining of the base model. Models left to right (blue dots): Pythia 70M, Pythia 160M, Pythia 410M, Pythia 1B, Pythia 2.8B, Pythia 6.9B, Llama 2 7B, Llama 3 8B Inst, Llama 3 70B Inst.
We find that mean task accuracy after fine-tuning improves with pretraining compute (Figure 7). However, we find signs of diminishing returns. For the Supreme Court tasks, performance increases steadily from $10^{20}$ to $10^{24}$ FLOPs, but further scaling to $10^{25}$ FLOPs only improves performance by an additional 4.0 accuracy points. For Appeals Court tasks, performance sharply increases from $10^{20}$ to $10^{21}$ FLOPs (i.e., Pythia 1B – which interestingly already beats GPT-4 zero-shot), but stagnates thereafter, only improving by an additional 8.5 accuracy points when scaling to $10^{25}$ FLOPs (i.e., Llama 3 70B Inst).
Our findings suggest that major improvements will likely not come from model scale alone. Rather, future work should focus on obtaining better scaling behavior. One promising direction is to improve the quality, quantity and diversity of the fine-tuning data.
3.2 Sample efficiency
We study how task accuracy scales as models fine-tune on more training examples. We consider the 10 tasks highlighted in Section 2.4. We fine-tune Llama 3 8B Instruct on each task independently, rather than on all tasks simultaneously as in the previous experiments. For each task, we fine-tune on 10, 50, 100, 250, 500, and 1000 task examples. We select task examples uniformly at random, and train 5 different models corresponding to different random seeds on the examples selected for training. We therefore fine-tune and evaluate a total of $10\cdot 6\cdot 5=300$ models. We fine-tune for a maximum of 20 epochs and early stop when validation loss increases for 3 consecutive epochs.
Figure 8: Sample efficiency of fine-tuning Llama 3 8B on a single task. Hundreds of task examples are typically enough to match or beat the zero-shot performance of GPT-4. Dashed blue line indicates the accuracy of Llama 3 8B fine-tuned on a single task as a function of the number of training examples. The shaded area indicates the 95% confidence interval over the randomly sampled training examples (5 random seeds)
Figure 8 shows how accuracy improves with the number of training examples. Fifty training examples are enough to match or beat the GPT-4 zero-shot baseline for 6 out of the 10 highlighted tasks, and 250 traning examples are enough to match or beat GPT-4 for 8 out of the 10 highlighted tasks. This is crucial, since labeling a few hundred data points is often financially feasible for many legal scholars (Hall and Wright, 2008). With relative few labelled task examples, fine-tuning reasonably small publicly available models can be competitive with state-of-the-art closed models. Moreover, accuracy continues to improve significantly with additional examples. With one thousand training examples, fine-tuning Llama 3 8B Inst matches or beats the GPT-4 baseline for all of the highlighted tasks.
3.3 Specializing for single tasks
We now study how much accuracy we stand to gain by fine-tuning on a single task. We specialize models for each of the 10 tasks highlighted in Section 2.4. We specialize the follow models: Llama 3 8B Inst, Llama 3 8B Inst fine-tuned for one epoch on all tasks, and Lawma 8B (i.e., Llama 3 8B Inst fine-tuned for three epochs on all tasks). For each task, we fine-tune for a maximum of 20 epochs and early stop when validation loss increases for 3 consecutive evaluation stpes, each corresponding to one tenth of an epoch.
Figure 9 shows the results of specialization to single tasks. First, we observe that, for 7 out of 10 tasks, Llama 3 8B Inst fine-tuned on all tasks for one epoch (yellow) outperforms Llama 3 8B Inst specialized for a single task (blue). That is, there is value to fine-tuning on our entire dataset rather than overspecializing for a single task. One explanation is that there is substantial cross-task overlap, and fine-tuning on the entire dataset amounts training on many more examples –even if on average these examples are less relevant.
Secondly, we observe that after fine-tuning on all 260 tasks for 1 epoch (yellow), further specializing for a single task (green) improves performance on all cases. Importantly, the latter outperforms the specialized Llama 3 8B Inst (blue) in all tasks. That is, a model that is fine-tuned on everything provides a “better” foundation from which to then “overspecializing” for a single task.
Thirdly fine-tuning on everything for three epochs (i.e., Lawma 8B, in red) again improves over the specialized models (i.e., green). Lastly, “overspecializing” Lawma 8B for a single task results in small single digit improvements for 3 out of the 10 tasks. However, we observe no benefits from specializing Lawma 8B for most (7/10) of the tasks. [^4] There is a small decrease in performance for SC Issue Area. This is because early stopping is performed with respect to loss on the validation set, but models are evaluated for accuracy on the test set. These results show that we don’t leave much accuracy on the table by fine-tuning a single model for all tasks. This is practically quite appealing, since it obviates the need to maintain a separate model for each task. A single model suffices.
Figure 9: Specializing Lawma 8B to individual tasks can yield small improvements in accuracy.
3.4 Generalization to unseen databases
We now investigate whether fine-tuning only on the Songer Appeals Court database allows us to generalize to the Supreme Court database. We fine-tune Llama 3 8B Inst for one epoch on all Songer tasks simultaneously. We plot in Figure 10 the mean accuracy for Court of Appeals tasks and Supreme Court tasks at intermediate checkpoints. As expected, performance on Court of Appeals tasks improves monotonically with the number of training examples seen. More interestingly, we observe that mean task accuracy for the Supreme Court also improves substantially, by up to 18.8 accuracy points at 20% of the training steps [^5] Note that 20% is the optimal amount of Songer data to train on if the goal is to generalize precisely to Supreme Court cases. If the goal is to generalize to some other dataset, then 20% need not be the optimal amount of Songer data to train on.. Thereafter, performance degrades, seemingly plateauing at 11.3 accuracy points above the baseline non-finetuned performance of Llama 3 8B Inst.
Our findings indicate that, since there is some degree of overlap between Court of Appeal and Supreme Court tasks, fine-tuning on the former transfers to the latter. This suggests that Lawma might be of practical use beyond the Supreme Court and Court of Appeals tasks it was trained on.
Note, however, that fine-tuning only on the Court of Appeals database results in a mean case accuracy of 51.6%, compared to 82.4% for Lawma 8B. That is, not fine-tuning on Supreme Court cases results in a 30.9 accuracy points decrease in performance. These results again highlight the importance of fine-tuning precisely on the target tasks of interest.
Figure 10: Fine-tuning only on the Court of Appeals database improves accuracy on Supreme Court database.
3.5 Intercoder agreement analysis
The Songer Appeals Court database provides intercoder agreement rates for a subset of the variables. These intercoder agreement rates provide valuable context for the performance of our model. Specifically, intercoder agreement gives us information about the inherent label noise in the annotation procedure. In particular, the intercoder agreement rate gives a natural upper bound on model performance, as we cannot expect the model to perform well when the label is uncertain or subject to interpretation.
However, we cannot directly compare intercoder agreement rates with the accuracy numbers we report. The reason is that in each task we subsampled the majority class to be no larger than the union of all other classes. This is a design choice we made to account for class imbalance. In this section, we map our model’s accuracy to adjusted accuracy numbers that undo the subsampling step. This results in accuracy numbers that are commensurate with the intercoder agreement rate.
Name IC Agreement Adj accuracy (unadjusted) Keep
WEIGHTEV (songer_weightev) 76 78.7% (77.2%) 28.72%
PROCEDUR (songer_procedur) 78 75.2% (73.9%) 83.08%
ORIGIN (songer_origin) 83.2 80.1% (77.7%) 53.13%
DIRECT2 (songer_direct2) 85.6 67.5% (67.5%) 100.00%
DIRECT1 (songer_direct1) 94 80.5% (80.5%) 100.00%
TREAT (songer_treat) 95.2 91.1% (90.1%) 71.26%
GENISS (songer_geniss) 97.6 93.2% (92.9%) 84.77%
CIRCUIT (songer_circuit) 100 93.2% (93.2%) 100.00%
COMMENT (songer_comment) 100 100.0% (91.7%) 0.13%
Table 2: Intercoder agreement rates, Lawma accuracies, and fraction of the majority class retained in our sample. Rows are sorted in increasing order of agreement rate.
Table 2 considers several tasks from the Appeals Court database, including the selected ones we highlighted in various figures. Each row corresponds to one task and provides the intercoder agreement rate, adjusted (and unadjusted) accuracy achieved by Lawma 8B, and the fraction of samples we retained in the majority class. A fraction of 100% means that we kept all samples. The smaller the fraction the larger the majority class is relative to the other classes.
The table contains several interesting insights:
  • The adjusted accuracy of Lawma 8B is generally within single digit percentage points of the intercoder agreement rate for easy tasks such as general issue classification (GENISS).
  • Lawma 8B is surprisingly close on the two tasks with the lowest intercoder reliability, i.e., WEIGHTEV and PROCEDUR. This shows that high intercoder reliability is no prerequesite for the model to perform well, i.e., close to the agreement rate.
  • On harder tasks, like identifying the ideological valence of a decision (DIRECT1 and DIRECT2), Lawma 8B is below the agreement rate by double digit percentage points.
  • Tasks with very high agreement rate (e.g., CIRCUIT and COMMENT) are not all alike. Some of them (e.g., COMMENT) correspond to a task with extreme class imbalance. Here, the model reaches the agreement rate. Other tasks (e.g., CIRCUIT) have perfect agreement rate, no class imbalance, and yet Lawma is far from the agreement rate.
These findings speak to the task heterogeneity and the non-trivial nature of the task suite as a classification benchmark.
4 Discussion
The cost of human annotators represents a considerable bottleneck for the field of empirical legal studies. In many scientific disciplines, the advent of low-cost and flexible tools for data extraction can lead to tremendous boosts in scholarly productivity and knowledge production. For example, the falling cost of genetic sequencing led to a paradigm shift across the biological sciences, as genetic data became increasingly available in fields as disparate as public health and entomology (Köser et al., 2012; Ballare et al., 2019). A flexible automated feature extraction tool for legal texts holds similar potential for empirical legal studies, as a large realm of conceivable but impracticably expensive research projects becomes accessible. In addition, such tools would boost the utility of existing legal databases.
The few-shot capabilities of large language models are vital for commercial APIs, where users are largely restricted to prompting. But as we show, zero-shot prompting is neither sufficiently good nor best possible for classification tasks that arise in empirical legal work. Lightly fine-tuned special purpose models achieve significantly higher accuracy from relatively few labeled examples. Labeling a few hundred cases is often financially feasible. This suggests a simple and practical strategy for solving legal classification tasks: Obtain a few hundred labeled examples, fine-tune an open source model, and use the fine-tuned model to annotate the remaining cases.
Multi-shot prompting is unlikely to provide much relief about the limitations of zero-shot prompting. Many legal documents are so long that a single document essentially consumes the entire prompt window. For example, 38% of court opinions in our dataset have more than 4,000 tokens. The default GPT-4 context size is 8,000 tokens. A version of GPT-4 allowing for 32,000 tokens is substantially more expensive. Fine-tuning avoids this problem altogether and robustly handles long documents.
The tasks we introduce are also interesting from a benchmarking perspective. The accuracy numbers are neither too low nor too high. The best models achieve non-trivial, but modest zero-shot performance. And even fine-tuned models don’t reach intercoder agreement rates. This situation suggests that these legal classification tasks may be good test cases for future model advances. As such, we hope to extend and strengthen existing evaluation efforts.
Acknowledgments
We would like to thank Nathan Adams, Jonathan Choi, James Grimmelmann, Neel Guha, Gillian Hadfield, Sanmi Koyejo, Katrina Ligett, Julian Nyarko, and Andrea Roth for helpful comments and feedback. R.A. was supported by the Andrew Carnegie Fellowship Program.
Appendix A Fine-tuning details
Compute requirements.
We fine-tune on a cluster consisting of NVIDIA H100 GPUs. Fine-tuning on all tasks simultaneously required approximately 600 H100 hours for the 8B model and 1600 GPU hours for the 70B model. In total, the experiments presented in the paper required approximately 8000 H100 GPU hours.
A.1 Lawma
We fine-tuning with a maximum sequence length of 8192 tokens. We use the AdamW optimizer with full precision, $\beta_{1}=0.9$, $\beta_{2}=0.95$, $\epsilon=10^{-8}$. We use a peak learning rate of $2\cdot 10^{-6}$. We use a cosine learning rate schedule, with 180 warm-up steps (approx. 4% of a full epoch) and decay to $10\%$ of the peak learning rate. We use a weight decay of 0.1. We clip gradient to 1.0 max norm. We pack samples using the axolotl library (Cloud, 2024), which improves training efficiency by approximately 40%. For Lawma 8B, we fine-tune Llama 3 8B Instruct for 3 epochs. We train on a node of 7 H100s using DeepSpeed Zero 2, with a global batch size of 56. For Lawma 70B, we fine-tune Llama 3 70B Instruct for 1 epoch. We train on 8 nodes of 8 H100s each using DeepSpeed Zero 3, with a global batch size of 64. We find that additional epochs hurt average task performance, although performance continues to improve for some of the tasks.
A.2 Additional fine-tuning experiments
The hyperparameters are identical to those used for Lawma unless otherwise specified.
Scaling experiments.
We fine-tune the Pythia and Llama 2 models with a peak learning rate of $2\cdot 10^{-5}$, which we find to be result in higher performance than a peak learning rate of $2\cdot 10^{-6}$. For the Llama 3 models, we use a learning rate of $2\cdot 10^{-6}$, which we find to be perform better than $2\cdot 10^{-5}$. We fine-tune for a single epoch. We use a batch size 64. We fine-tune models with their pretraining max sequence length, that is, 2k tokens for Pythia, 4k tokens for Llama 2, and 8k tokens for Llama 3. We use a warm up ratio of 0.03. Due to the costs associated with training the 70B model, we simply take Lawma 70B rather than re-training the model with these slightly different training hyperparameters.
Sample efficiency and specialization
We fine-tune for up to 20 epochs. We evaluate the loss on a separate validation set and early stop if the loss increases for 3 consecutive evaluation steps. For the sample efficiency experiments, we evaluate at the end of every epoch. For the specialization experiments, we evaluate every 0.1 epochs. We decay the learning rate to 10% of the peak learning rate over the 20 epochs. We fine-tune with a batch size of 64. For the specialization experiments, we train models both with and without learning rate warm up, and report the accuracy of the best model. We use the AdamW BitsAndBytes 8-bit optimizer, allowing us to fine-tune the models in a single H100 GPU.
Generalization
We fine-tune only on the Songer Court of Appeals tasks. We fine-tune with batch size 64. We fine-tune for one epoch and we checkpoint models at 10, 30, 60, 100, 300, 600, 1000, 2000, and 3000 training steps. A full epoch on the Songer Court of Appeal tasks corresponds to 3096 training steps.
Appendix B List of all tasks
Variable Question Sample answer choices
sc_adminaction What is the agency involved in the administrative action? Army and Air Force Exchange Service, Atomic Energy Commission, Secretary or administrative unit or personnel of the U.S. Air Force
sc_adminaction_is Did administrative action occur in the context of the case? No, Yes
sc_adminactionstate What is the state of the state agency associated with the administrative action? Alabama, Alaska, American Samoa
sc_authoritydecision What is the basis of the Supreme Court’s decision? judicial review (national level), judicial review (state level), Supreme Court supervision of lower federal or state courts or original jurisdiction
sc_casedisposition What is the disposition of the case, that is, the treatment the Supreme Court accorded the court whose decision it reviewed? stay, petition, or motion granted, affirmed (includes modified), reversed
sc_caseorigin What is the court in which the case originated? U.S. Court of Customs and Patent Appeals, U.S. Court of International Trade, U.S. Court of Claims, Court of Federal Claims
sc_caseoriginstate What is the state of the court in which the case originated? Alabama, Alaska, American Samoa
sc_casesource What is the court whose decision the Supreme Court reviewed? U.S. Court of Customs and Patent Appeals, U.S. Court of International Trade, U.S. Court of Claims, Court of Federal Claims
sc_casesourcestate What is the state of the court whose decision the Supreme Court reviewed? Alabama, Alaska, American Samoa
sc_certreason What reason, if any, does the court give for granting the petition for certiorari? case did not arise on cert or cert not granted, federal court conflict, federal court conflict and to resolve important or significant question
sc_decisiondirection What is the ideological direction of the decision? Conservative, Liberal, Unspecifiable
sc_decisiontype What type of decision did the court make? opinion of the court (orally argued), per curiam (no oral argument), decrees
sc_declarationuncon Did the Court declare unconstitutional an act of Congress; a state or territorial statute, regulation, or constitutional provision; or a municipal or other local ordinance? No declaration of unconstitutionality, Act of Congress declared unconstitutional, State or territorial law, regulation, or constitutional provision unconstitutional
sc_issue_1 What is the issue of the decision? subconstitutional fair procedure: fugitive from justice, self-incrimination, immunity from prosecution, cruel and unusual punishment, death penalty (cf. extra legal jury influence, death penalty)
sc_issue_10 What is the issue of the decision? federal pre-emption of state legislation or regulation. cf. state regulation of business. rarely involves union activity. Does not involve constitutional interpretation unless the Court says it does., federal pre-emption of state legislation or regulation. cf. state regulation of business. rarely involves union activity. Does not involve constitutional interpretation unless the Court says it does., national supremacy: public utilities (cf. federal public utilities regulation)
sc_issue_11 What is the issue of the decision? non-real property dispute between states, non-real property dispute between states, boundary dispute between states
sc_issue_12 What is the issue of the decision? federal taxation, typically under provisions of the Internal Revenue Code, federal taxation, typically under provisions of the Internal Revenue Code, federal taxation of gifts, personal, business, or professional expenses
sc_issue_2 What is the issue of the decision? sex discrimination (excluding sex discrimination in employment), Voting Rights Act of 1965, plus amendments, juveniles (cf. rights of illegitimates)
sc_issue_3 What is the issue of the decision? libel, privacy: true and false light invasions of privacy, parochiaid: government aid to religious schools, or religious requirements in public schools, First Amendment, miscellaneous (cf. comity: First Amendment)
sc_issue_4 What is the issue of the decision? due process: takings clause, or other non-constitutional governmental taking of property, due process: miscellaneous (cf. loyalty oath), the residual code, due process: miscellaneous (cf. loyalty oath), the residual code
sc_issue_5 What is the issue of the decision? Freedom of Information Act and related federal or state statutes or regulations, abortion: including contraceptives, abortion: including contraceptives
sc_issue_6 What is the issue of the decision? attorneys’ and governmental employees’ or officials’ fees or compensation or licenses, commercial speech, attorneys (cf. commercial speech), attorneys’ and governmental employees’ or officials’ fees or compensation or licenses
sc_issue_7 What is the issue of the decision? labor-management disputes: right to organize, union-union member dispute (except as pertains to union or closed shop), labor-management disputes: employee discharge
sc_issue_8 What is the issue of the decision? natural resources - environmental protection (cf. national supremacy: natural resources, national supremacy: pollution), Employee Retirement Income Security Act (cf. union trust funds), election of remedies: legal remedies available to injured persons or things
sc_issue_9 What is the issue of the decision? standing to sue: private or implied cause of action, judicial administration: review of non-final order, judicial administration: jurisdiction or authority of federal district courts or territorial courts
sc_issuearea What is the issue area of the decision? Criminal Procedure, Civil Rights, First Amendment
sc_jurisdiction What is the manner in which the Court took jurisdiction? cert, appeal, bail
sc_lcdisagreement Does the court opinion mention that one or more of the members of the court whose decision the Supreme Court reviewed dissented? Yes, No
sc_lcdisposition What treatment did the court whose decision the Supreme Court reviewed accorded the decision of the court it reviewed? stay, petition, or motion granted, affirmed, reversed
sc_lcdispositiondirection What is the ideological direction of the decision reviewed by the Supreme Court? Conservative, Liberal, Unspecifiable
sc_partywinning Consider that the petitioning party lost if the Supreme Court affirmed or dismissed the case, or denied the petition. Consider that the petitioning party won in part or in full if the Supreme Court reversed, reversed and remanded, vacated and remanded, affirmed and reversed in part, affirmed and reversed in part and remanded, or vacated the case. Did the petitioning win the case? Yes, No
sc_petitioner Who is the petitioner of the case? attorney general of the United States, or his office, specified state board or department of education, city, town, township, village, or borough government or governmental unit
sc_petitionerstate What state is associated with the petitioner? Alabama, Alaska, American Samoa
sc_precedentalteration Did the the decision of the court overrule one or more of the Court’s own precedents? Yes, No
sc_respondent Who is the respondent of the case? attorney general of the United States, or his office, specified state board or department of education, city, town, township, village, or borough government or governmental unit
sc_respondentstate What state is associated with the respondent? Alabama, Alaska, American Samoa
sc_threejudgefdc Was the case heard by a three-judge federal district court? Yes, No
songer_abusedis Did the court conclude that it should defer to agency discretion? For example, if the action was committed to agency discretion. No, Yes, Mixed answer
songer_adminrev What federal agency’s decision was reviewed by the court of appeals? Benefits Review Board, Civil Aeronautics Board, Civil Service Commission
songer_agen_acq Did the court rule for the government in an issue related to agency acquisition of information (e.g. physical inspections, searches, subpoenas, records, etc)? No, Yes, Mixed answer
songer_alj Did the court support the decision of an administrative law judge? No, Yes, Mixed answer
songer_altdisp Did the court’s ruling on an issue arising out of an alternative dispute resolution process (ADR, settlement conference, role of mediator or arbitrator, etc.) favor the appellant? No, Yes, Mixed answer
songer_amicus Was there any amicus participation before the court of appeals? no amicus participation on either side, 1 separate amicus brief was filed, 2 separate amicus briefs were filed
songer_app_stid What is the state of the first listed state or local government agency that is an appellant? not, Alabama, Alaska
songer_appbus What is the total number of appellants in the case that fall into the category "private business and its executives"? Answer with a number. N/A
songer_appel1_1_2 This question concerns the first listed appellant. The nature of this litigant falls into the category "private business (including criminal enterprises)". What is the scope of this business? local, neither local nor national, national or multi-national
songer_appel1_1_3 This question concerns the first listed appellant. The nature of this litigant falls into the category "private business (including criminal enterprises)". What category of business best describes the area of activity of this litigant which is involved in this case? agriculture, mining, construction
songer_appel1_1_4 This question concerns the first listed appellant. The nature of this litigant falls into the category "private business (including criminal enterprises)", specifically "agriculture". What subcategory of business best describes this litigant? single family farm, commercial farm, agri-business, farm - other
songer_appel1_2_2 This question concerns the first listed appellant. The nature of this litigant falls into the category "private organization or association". What category of private associations best describes this litigant? business, trade, professional, or union (BTPU), other
songer_appel1_2_3 This question concerns the first listed appellant. The nature of this litigant falls into the category "private organization or association", specifically "business, trade, professional, or union (BTPU)". What subcategory of private association best describes this litigant? Business or trade association, utilities co-ops, Professional association - other than law or medicine
songer_appel1_3_2 This question concerns the first listed appellant. The nature of this litigant falls into the category "federal government (including DC)". Which category of federal government agencies and activities best describes this litigant? cabinet level department, courts or legislative, agency whose first word is "federal"
songer_appel1_3_3 This question concerns the first listed appellant. The nature of this litigant falls into the category "federal government (including DC)", specifically "cabinet level department". Which specific federal government agency best describes this litigant? Department of Agriculture, Department of Commerce, Department of Defense (includes War Department and Navy Department)
songer_appel1_4_2 This question concerns the first listed appellant. The nature of this litigant falls into the category "sub-state government (e.g., county, local, special district)". Which category of substate government best describes this litigant? legislative, executive/administrative, bureaucracy providing services
songer_appel1_4_3 This question concerns the first listed appellant. The nature of this litigant falls into the category "sub-state government (e.g., county, local, special district)", specifically "legislative". Which specific substate government agency best describes this litigant? City/county council, School Board, board of trustees for college or junior college, Other legislative body
songer_appel1_5_2 This question concerns the first listed appellant. The nature of this litigant falls into the category "state government (includes territories & commonwealths)". Which category of state government best describes this litigant? legislative, executive/administrative, bureaucracy providing services
songer_appel1_5_3 This question concerns the first listed appellant. The nature of this litigant falls into the category "state government (includes territories & commonwealths)", specifically "legislative". Which specific state government agency best describes this litigant? Legislature or separate house as an organization, Legislative Committee or Commission, Other Legislative Unit
songer_appel1_7_2 This question concerns the first listed appellant. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the gender of this litigant?Use names to classify the party’s sex only if there is little ambiguity. not ascertained, male - indication in opinion (e.g., use of masculine pronoun), male - assumed because of name
songer_appel1_7_3 This question concerns the first listed appellant. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the race or ethnic identity of this litigant as identified in the opinion? not ascertained, caucasian - specific indication in opinion, black - specific indication in opinion
songer_appel1_7_4 This question concerns the first listed appellant. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the citizenship of this litigant as indicated in the opinion? not ascertained, US citizen, alien
songer_appel1_7_5 This question concerns the first listed appellant. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". Which of these categories best describes the income of the litigant? not ascertained, poor + wards of state, presumed poor
songer_appel1_8_2 This question concerns the first listed appellant. The nature of this litigant falls into the category "miscellaneous". Which of the following categories best describes the litigant? fiduciary, executor, or trustee, other, nature of the litigant not ascertained
songer_appel1_8_3 This question concerns the first listed appellant. The nature of this litigant falls into the category "miscellaneous", specifically "fiduciary, executor, or trustee". Which of the following specific subcategories best describes the litigant? trustee in bankruptcy - institution, trustee in bankruptcy - individual, executor or administrator of estate - institution
songer_appel2_1_2 This question concerns the second listed appellant. The nature of this litigant falls into the category "private business (including criminal enterprises)". What is the scope of this business? local, neither local nor national, national or multi-national
songer_appel2_1_3 This question concerns the second listed appellant. The nature of this litigant falls into the category "private business (including criminal enterprises)". What category of business best describes the area of activity of this litigant which is involved in this case? agriculture, mining, construction
songer_appel2_1_4 This question concerns the second listed appellant. The nature of this litigant falls into the category "private business (including criminal enterprises)", specifically "agriculture". What subcategory of business best describes this litigant? single family farm, commercial farm, agri-business, farm - other
songer_appel2_2_2 This question concerns the second listed appellant. The nature of this litigant falls into the category "private organization or association". What category of private associations best describes this litigant? business, trade, professional, or union (BTPU), other
songer_appel2_2_3 This question concerns the second listed appellant. The nature of this litigant falls into the category "private organization or association", specifically "business, trade, professional, or union (BTPU)". What subcategory of private association best describes this litigant? Business or trade association, utilities co-ops, Professional association - other than law or medicine
songer_appel2_3_2 This question concerns the second listed appellant. The nature of this litigant falls into the category "federal government (including DC)". Which category of federal government agencies and activities best describes this litigant? cabinet level department, courts or legislative, agency whose first word is "federal"
songer_appel2_3_3 This question concerns the second listed appellant. The nature of this litigant falls into the category "federal government (including DC)", specifically "cabinet level department". Which specific federal government agency best describes this litigant? Department of Agriculture, Department of Commerce, Department of Defense (includes War Department and Navy Department)
songer_appel2_4_2 This question concerns the second listed appellant. The nature of this litigant falls into the category "sub-state government (e.g., county, local, special district)". Which category of substate government best describes this litigant? legislative, executive/administrative, bureaucracy providing services
songer_appel2_4_3 This question concerns the second listed appellant. The nature of this litigant falls into the category "sub-state government (e.g., county, local, special district)", specifically "legislative". Which specific substate government agency best describes this litigant? City/county council, School Board, board of trustees for college or junior college, Other legislative body
songer_appel2_5_2 This question concerns the second listed appellant. The nature of this litigant falls into the category "state government (includes territories & commonwealths)". Which category of state government best describes this litigant? legislative, executive/administrative, bureaucracy providing services
songer_appel2_5_3 This question concerns the second listed appellant. The nature of this litigant falls into the category "state government (includes territories & commonwealths)", specifically "legislative". Which specific state government agency best describes this litigant? Legislature or separate house as an organization, Legislative Committee or Commission, Other Legislative Unit
songer_appel2_7_2 This question concerns the second listed appellant. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the gender of this litigant?Use names to classify the party’s sex only if there is little ambiguity. not ascertained, male - indication in opinion (e.g., use of masculine pronoun), male - assumed because of name
songer_appel2_7_3 This question concerns the second listed appellant. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the race or ethnic identity of this litigant as identified in the opinion? not ascertained, caucasian - specific indication in opinion, black - specific indication in opinion
songer_appel2_7_4 This question concerns the second listed appellant. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the citizenship of this litigant as indicated in the opinion? not ascertained, US citizen, alien
songer_appel2_7_5 This question concerns the second listed appellant. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". Which of these categories best describes the income of the litigant? not ascertained, poor + wards of state, presumed poor
songer_appel2_8_2 This question concerns the second listed appellant. The nature of this litigant falls into the category "miscellaneous". Which of the following categories best describes the litigant? fiduciary, executor, or trustee, other, nature of the litigant not ascertained
songer_appel2_8_3 This question concerns the second listed appellant. The nature of this litigant falls into the category "miscellaneous", specifically "fiduciary, executor, or trustee". Which of the following specific subcategories best describes the litigant? trustee in bankruptcy - institution, trustee in bankruptcy - individual, executor or administrator of estate - institution
songer_appfed What is the total number of appellants in the case that fall into the category "the federal government, its agencies, and officialss"? Answer with a number. N/A
songer_appfiduc What is the total number of appellants in the case that fall into the category "fiduciaries"? Answer with a number. N/A
songer_applfrom What is the type of district court decision or judgment appealed from (i.e., the nature of the decision below in the district court)? Trial (either jury or bench trial), Injunction or denial of injunction or stay of injunction, Summary judgment or denial of summary judgment
songer_appnatpr What is the total number of appellants in the case that fall into the category "natural persons"? Answer with a number. N/A
songer_appnonp What is the total number of appellants in the case that fall into the category "groups and associations"? Answer with a number. N/A
songer_appstate What is the total number of appellants in the case that fall into the category "state governments, their agencies, and officials"? Answer with a number. N/A
songer_appsubst What is the total number of appellants in the case that fall into the category "sub-state governments, their agencies, and officials"? Answer with a number. N/A
songer_attyfee Did the court’s ruling on attorneys’ fees favor the appellant? No, Yes, Mixed answer
songer_bank_app1 Is the first listed appellant bankrupt? Yes, No
songer_bank_app2 Is the second listed appellant bankrupt? Yes, No
songer_bank_r1 Is the first listed respondent bankrupt? Yes, No
songer_bank_r2 Is the second listed respondent bankrupt? Yes, No
songer_capric Did the courts’s use or interpretation of the arbitrary and capricious standard support the government? Note that APA allows courts to overturn agency actions deemed to be arbitrary or capricious, an abuse of discretion, or otherwise not in accordance with law. Overton Park emphasized this is a narrow standard, and one must prove that agency’s action is without a rational basis. This also includes the "substantial justification" doctrine. No, Yes, Mixed answer
songer_casetyp1_1-2 What is the specific issue in the case within the general category of "issue"? federal offense, state offense, not determined whether state or federal offense
songer_casetyp1_1-3-1 What is the specific issue in the case within the general category of "issue"? murder, rape, arson
songer_casetyp1_1-3-2 What is the specific issue in the case within the general category of "issue"? murder, rape, arson
songer_casetyp1_1-3-3 What is the specific issue in the case within the general category of "issue"? murder, rape, arson
songer_casetyp1_2-2 What is the specific issue in the case within the general category of "issue"? civil rights claims by prisoners and those accused of crimes, voting rights, race discrimination, sex discrimination, other civil rights
songer_casetyp1_2-3-1 What is the specific issue in the case within the general category of "issue"? suit for damages for false arrest or false confinement, cruel and unusual punishment, due process rights in prison
songer_casetyp1_2-3-2 What is the specific issue in the case within the general category of "issue"? voting rights - reapportionment & districting, participation rights - rights of candidates or groups to fully participate in the political process; access to ballot, voting rights - other (includes race discrimination in voting)
songer_casetyp1_2-3-3 What is the specific issue in the case within the general category of "issue"? alien petitions - (includes disputes over attempts at deportation), indian rights and law, juveniles
songer_casetyp1_3-2 What is the specific issue in the case within the general category of "issue"? religion, press, commercial, speech and other expression
songer_casetyp1_3-3-1 What is the specific issue in the case within the general category of "issue"? commercial speech, libel, slander, defamation, free exercise of religion
songer_casetyp1_3-3-2 What is the specific issue in the case within the general category of "issue"? obscenity, association, federal internal security and communist control acts, loyalty oaths, security risks
songer_casetyp1_4-3 What is the specific issue in the case within the general category of "issue"? denial of fair hearing or notice - government employees (includes claims of terminated government workers), denial of hearing or notice in non-employment context, taking clause (i.e., denial of due process under the "taking" clause of the 5th or 14th Amendments)
songer_casetyp1_5-3 What is the specific issue in the case within the general category of "issue"? abortion rights, homosexual rights where privacy claim raised, contraception and other privacy claims related to marital relations or sexual behavior (not in 501 or 502)
songer_casetyp1_6-3 What is the specific issue in the case within the general category of "issue"? union organizing, unfair labor practices, Fair Labor Standards Act issues
songer_casetyp1_7-2 What is the specific issue in the case within the general category of "issue"? taxes, patents, copyright, torts, commercial disputes
songer_casetyp1_7-3-1 What is the specific issue in the case within the general category of "issue"? state or local tax, federal taxation - individual income tax (includes taxes of individuals, fiduciaries, & estates), federal tax - business income tax (includes corporate and parnership)
songer_casetyp1_7-3-2 What is the specific issue in the case within the general category of "issue"? motor vehicle, airplane, product liability
songer_casetyp1_7-3-3 What is the specific issue in the case within the general category of "issue"? contract disputes-general (private parties) (includes breach of contract, disputes over meaning of contracts, suits for specific performance, disputes over whether contract fulfilled, claims that money owed on contract) (Note: this category is not used when the dispute fits one of the more specific categories below), disputes over government contracts, insurance disputes
songer_casetyp1_7-3-4 What is the specific issue in the case within the general category of "issue"? bankruptcy - private individual (e.g., chapter 7), bankruptcy - business reorganization (e.g., chapter 11), other bankruptcy
songer_casetyp1_7-3-5 What is the specific issue in the case within the general category of "issue"? social security benefits (including SS disability payments), other government benefit programs (e.g., welfare, RR retirement, veterans benefits, war risk insurance, food stamps), state or local economic regulation
songer_casetyp1_7-3-6 What is the specific issue in the case within the general category of "issue"? disputes over real property (private), eminent domain and disputes with government over real property, landlord - tenant disputes
songer_casetyp1_9-3 What is the specific issue in the case within the general category of "issue"? miscellaneous interstate conflict, other federalism issue (only code as issue if opinion explicitly discusses federalism as an important issue - or if opinion explicity discusses conflict of state power vs federal power), attorneys (disbarment; etc)
songer_casetyp2_geniss What is the second general issue in the case, other than mainissue? criminal, civil rights, First Amendment
songer_circuit What is the circuit of the court that decided the case? First Circuit, Second Circuit, Third Circuit
songer_civproc1 What is the most frequently cited federal rule of civil procedure in the headnotes to this case? Answer with a number. N/A
songer_civproc2 What is the second most frequently cited federal rule of civil procedure in the headnotes to this case? Answer with a number. N/A
songer_classact Is the case described in the opinion as a class action suit? No, Yes
songer_comment Did the agency give proper opportunity to comment? No, Yes, Mixed answer
songer_concur What is the number of judges who concurred in the result but not in the opinion of the court? 0, 1, 2
songer_confess Did the court conclude that a confession or an incriminating statement was improperly admitted? Consider only incriminating statements made by the defendant. No, Yes, Yes, but error was harmless
songer_const1 What is the most frequently cited provision of the U.S. Constitution in the headnotes to this case? If it is one of the original articles of the constitution, code the number of the article preceeded by two zeros. If it is an amendment to the constitution, code the number of the amendment (zero filled to two places) preceeded by a "1". Examples: 001 = Article 1 of the original constitution, 101 = 1st Amendment, 114 = 14th Amendment. N/A
songer_const2 What is the second most frequently cited provision of the U.S. Constitution in the headnotes to this case? If it is one of the original articles of the constitution, code the number of the article preceeded by two zeros. If it is an amendment to the constitution, code the number of the amendment (zero filled to two places) preceeded by a "1". Examples: 001 = Article 1 of the original constitution, 101 = 1st Amendment, 114 = 14th Amendment. N/A
songer_constit Did the court’s conclusion about the constitutionality of a law or administrative action favor the appellant? Issue not discussed, The issue was discussed in the opinion and the resolution of the issue by the court favored the respondent, The issue was discussed in the opinion and the resolution of the issue by the court favored the appellant
songer_counsel Did the court rule that the defendant had inadequate counsel? No, Yes, Yes, but error was harmless
songer_counsel1 What is the nature of the counsel for the appellant? none (pro se), court appointed, legal aid or public defender
songer_counsel2 What is the nature of the counsel for the respondent? none (pro se), court appointed, legal aid or public defender
songer_crmproc1 What is the most frequently cited federal rule of criminal procedure in the headnotes to this case? Answer with a number. N/A
songer_crmproc2 What is the second most frequently cited federal rule of criminal procedure in the headnotes to this case? Answer with a number. N/A
songer_crossapp Were there cross appeals from the decision below to the court of appeals that were consolidated in the present case? No, Yes, Not ascertained
songer_deathpen Did the court conclude that the death penalty was improperly imposed? Consider only the validity of the sentence, rather than whether or not the conviction was proper. No, Yes, Yes, but error was harmless
songer_decuncon Did the court declare any statute or administrative action unconstitutional? no declarations of unconstitutionality, act of Congress declared unconstitutional (facial invalidity), interpretation/application of federal law invalid
songer_denovo Did the court’s use of the standard of review, "de novo on facts" support the government? The courts generally recognize that de novo review is impractical for the bulk of agency decisions so the substantial evidence standard helps provide a middle course. Consider the de novo review of administrative action, not de novo review of trial court by appeals court. No, Yes, Mixed answer
songer_direct1 What is the ideological directionality of the court of appeals decision? conservative, liberal, mixed
songer_direct2 What is the ideological directionality of the court of appeals decision? conservative, liberal, mixed
songer_discover Did the court’s interpretation of rules relating to discovery or other issues related to obtaining evidence favor the appellant? No, Yes, Mixed answer
songer_dissent What is the number of judges who dissented from the majority? 0, 1, 2
songer_district From which district in the state was this case appealed? Not applicable, Eastern, Western
songer_diverse Did the court conclude that the parties were truly diverse? No, Yes, Mixed answer
songer_dueproc Did the interpretation of the requirements of due process by the court favor the appellant? No, Yes, Mixed answer
songer_entrap Did the court rule that the defendant was the victim of illegal entrapment? No, Yes, Yes, but error was harmless
songer_erron Did the court’s use of the clearly erroneous standard support the government? That is, a somewhat narrower standard than substantial evidence, or ignoring usual agency standards. No, Yes, Mixed answer
songer_execord Did the interpretation of executive order or administrative regulation by the court favor the appellant? This does include whether or not an executive order was lawful. No, Yes, Mixed answer
songer_exhaust Did the court determine that it would not hear the appeal for one of the following reasons: a) administrative remedies had not been exhausted; or b) the issue was not ripe for judicial action? No, Yes, Mixed answer
songer_fedlaw Did the interpretation of federal statute by the court favor the appellant? No, Yes, Mixed answer
songer_fedvst Did the court rule that federal law should take precedence over state or local laws in a case involving the conflict of laws (i.e, which laws or rules apply)? No, Yes, Mixed answer
songer_foreign Did the court rule that domestic law (federal, state or local) should take precedence over foreign law in a case involving the conflict of laws (i.e., which laws or rules apply- foreign country vs federal, state, or local)? No, Yes, Mixed answer
songer_freeinfo Did the court rule in favor of the government when the administrative action in question related to the agency’s providing information to those who request it? For example, Freedom of Information, issues of governmental confidentiality, or "government in the sunshine". No, Yes, Mixed answer
songer_frivapp Did the court conclude that it could not reach the merits of the case because the motion or appeal was frivolous or raised only trivial issues and was therefore not suitable for appellate review? No, Yes, Mixed answer
songer_frivol Did the court conclude that either the original case was frivolous or raised only trivial issues and therefore was not suitable for actions on the merits? No, Yes, Mixed answer
songer_genapel1 What is the nature of the first listed appellant? private business (including criminal enterprises), private organization or association, federal government (including DC)
songer_genapel2 What is the nature of the second listed appellant whose detailed code is not identical to the code for the first listed appellant? private business (including criminal enterprises), private organization or association, federal government (including DC)
songer_geniss What is the general issue in the case? criminal, civil rights, First Amendment
songer_genresp1 What is the nature of the first listed respondent? private business (including criminal enterprises), private organization or association, federal government (including DC)
songer_genresp2 What is the nature of the second listed respondent whose detailed code is not identical to the code for the first listed respondent? private business (including criminal enterprises), private organization or association, federal government (including DC)
songer_genstand Did the agency articulate the appropriate general standard? This question includes whether the agency interpreted the statute "correctly". The courts often refer here to the rational basis test, plain meaning, reasonable construction of the statute, congressional intent, etc. This issue also includes question of which law applies or whether amended law vs law before amendment applies. No, Yes, Mixed answer
songer_habeas Was the case an appeal of a decision by the district court on a petition for habeas corpus? no, yes, state habeas corpus (criminal), yes, federal habeas corpus (criminal)
songer_immunity Did the court refuse to reach the merits of the appeal because it concluded that the defendant had immunity? No, Yes, Mixed answer
songer_improper Did the court conclude that there was improper influence on the jury? For example, include jury tampering or failure to shield jury from prejudicial media accounts. Exclude prejudicial conduct by the prosecutor. No, Yes, Yes, but error was harmless
songer_indict Did the court rule that the indictment was defective? No, Yes, Yes, but error was harmless
songer_indigent Did the court rule that the defendant’s rights as an indigent were violated? No, Yes, Yes, but error was harmless
songer_initiate What party initiated the appeal? Original plaintiff, Original defendant, Federal agency representing plaintiff
songer_injunct Did the court’s ruling on the validity of an injunction or the denial of an injunction or a stay of injunction favor the appellant? No, Yes, Mixed answer
songer_insane Did the court below err in not permitting an insanity defense? No, Yes, Yes, but error was harmless
songer_int_law Did the court rule in favor of the appellant on an issue related to the interpretation of a treaty or international law? No, Yes, Mixed answer
songer_interven Did one or more individuals or groups seek to formally intervene in the appeals court consideration of the case? no intervenor in case, intervenor = appellant, intervenor = respondent
songer_judgdisc Did the court’s ruling on the abuse of discretion by the trial judge favor the appellant? This includes the issue of whether the judge actually had the authority for the action taken, but does not include questions of discretion of administrative law judges. No, Yes, Mixed answer
songer_judrev Did the court conclude the decision was subject to judicial review? While questions of fact are subject to limited review, questions of law are subject to full review. The problem becomes determining which are clear questions of law or fact as they are often "mixed". No, Yes, Mixed answer
songer_jurisdiction Did the court determine that it had jurisdiction to hear this case? No, Yes, Mixed answer
songer_juryinst Did the court conclude that the jury instructions were improper? No, Yes, Yes, but error was harmless
songer_late Did the court refuse to decide the appeal because the appellant failed to comply with some rule relating to timeliness of the appeal? No, Yes, Mixed answer
songer_majvotes What is the number of judges who voted in favor of the disposition favored by the majority? 0, 1, 2
songer_method What is the nature of the proceeding in the court of appeals for this case? decided by panel for first time (no indication of re-hearing or remand), decided by panel after re-hearing (second time this case has been heard by this same panel), decided by panel after remand from Supreme Court
songer_mootness Did the court conclude that an issue was moot? No, Yes, Mixed answer
songer_notice Decisions that affect life, liberty, or property must be preceded by adequate notice and an opportunity for a fair hearing. Did the agency give proper notice? No, Yes, Mixed answer
songer_numappel What is the total number of appellants in the case? Answer with a number. N/A
songer_numresp What is the total number of respondents in the case? Answer with a number. N/A
songer_opinstat Is the opinion writer identified in the opinion, or was the opinion per curiam? Signed, with reasons, Per curiam, with reasons, Not ascertained
songer_origin What type of court made the original decision? Federal district court (single judge), 3 judge district court, State court
songer_othadmis Did the court rule that some evidence, other than a confession made by the defendant or illegal search and seizure, was inadmissibile (or did ruling on appropriateness of evidentary hearing benefit the defendant)? No, Yes, Yes, but error was harmless
songer_othappth Did the court refuse to rule on the merits of the appeal because of some threshhold issue other than timeliness or frivolousness that was relevant on appeal but not at the original trial? No, Yes, Mixed answer
songer_othcrim Did the court rule for the defendant on grounds other than procedural grounds? For example, right to speedy trial, double jeopardy, confrontation, retroactivity, self defense. This includes the question of whether the defendant waived the right to raise some claim. No, Yes, Yes, but error was harmless
songer_othjury Did the court conclude that the jury composition or selection was invalid or that the jury was biased or tampered with? No, Yes, Yes, but error was harmless
songer_oththres Did the court refuse to rule on the merits of the appeal because of a threshhold issue other than lack of jurisdiction, standing, mootness, failure to state a claim, exhaustion, timeliness, immunity, frivolousness, or nonjusticiable political question? No, Yes, Mixed answer
songer_plea Did the court rule for the defendant on an issue related to plea bargaining? Plea bargain includes all challenges to plea. No, Yes, Yes, but error was harmless
songer_polquest Did the court refuse to rule on the merits of the case because it was considered to be a nonjusticiable "political question"? No, Yes, Mixed answer
songer_post_trl Did the court’s ruling on some post-trial procedure or motion (e.g., allocating court costs or post award relief) favor the appellant? This doe not include attorneys’ fees, but does include motions to set aside a jury verdict. No, Yes, Mixed answer
songer_prejud Was there prejudicial conduct by prosecution? No, Yes, Yes, but error was harmless
songer_pretrial Did the court’s rulings on pre-trial procedure favor the appellant? This includes whether or not there is a right to jury trial, whether the case should be certified as a class action, or whether a prospective party has a right to intervene in the case, but does not include rulings on motions for summary judgment. No, Yes, Mixed answer
songer_procdis Did the court uphold the dismissal by district court on procedural grounds? No, Yes, Yes, but error was harmless
songer_procedur Did the interpretation of federal rule of procedures, judicial doctrine, or case law by the court favor the appellant? No, Yes, Mixed answer
songer_r_bus What is the total number of respondents in the case that fall into the category "private business and its executives"? Answer with a number. N/A
songer_r_fed What is the total number of respondents in the case that fall into the category "the federal government, its agencies, and officialss"? Answer with a number. N/A
songer_r_fiduc What is the total number of respondents in the case that fall into the category "fiduciaries"? Answer with a number. N/A
songer_r_natpr What is the total number of respondents in the case that fall into the category "natural persons"? Answer with a number. N/A
songer_r_nonp What is the total number of respondents in the case that fall into the category "groups and associations"? Answer with a number. N/A
songer_r_state What is the total number of respondents in the case that fall into the category "state governments, their agencies, and officials"? Answer with a number. N/A
songer_r_stid What is the state of the first listed state or local government agency that is a respondent? not, Alabama, Alaska
songer_r_subst What is the total number of respondents in the case that fall into the category "sub-state governments, their agencies, and officials"? Answer with a number. N/A
songer_realapp Are the formally listed appellants in the case the "real parties", that is, are they the parties whose real interests are most directly at stake? both 1st and 2nd listed appellants are real parties (or only one appellant, and that appellant is a real party), the 1st appellant is not a real party, the 2nd appellant is not a real party
songer_realresp Are the formally listed respondents in the case the "real parties", that is, are they the parties whose real interests are most directly at stake? both 1st and 2nd listed respondents are real parties (or only one respondent, and that respondent is a real party), the 1st respondent is not a real party, the 2nd respondent is not a real party
songer_record Did the agency fail to develop an adequate record? For example, if the court was unable to determine what doctrine was used for the decision or unable to determine the basis of the decision. No, Yes, Mixed answer
songer_respond1_1_2 This question concerns the first listed respondent. The nature of this litigant falls into the category "private business (including criminal enterprises)". What is the scope of this business? local, neither local nor national, national or multi-national
songer_respond1_1_3 This question concerns the first listed respondent. The nature of this litigant falls into the category "private business (including criminal enterprises)". What category of business best describes the area of activity of this litigant which is involved in this case? agriculture, mining, construction
songer_respond1_1_4 This question concerns the first listed respondent. The nature of this litigant falls into the category "private business (including criminal enterprises)", specifically "agriculture". What subcategory of business best describes this litigant? single family farm, commercial farm, agri-business, farm - other
songer_respond1_2_2 This question concerns the first listed respondent. The nature of this litigant falls into the category "private organization or association". What category of private associations best describes this litigant? business, trade, professional, or union (BTPU), other
songer_respond1_2_3 This question concerns the first listed respondent. The nature of this litigant falls into the category "private organization or association", specifically "business, trade, professional, or union (BTPU)". What subcategory of private association best describes this litigant? Business or trade association, utilities co-ops, Professional association - other than law or medicine
songer_respond1_3_2 This question concerns the first listed respondent. The nature of this litigant falls into the category "federal government (including DC)". Which category of federal government agencies and activities best describes this litigant? cabinet level department, courts or legislative, agency whose first word is "federal"
songer_respond1_3_3 This question concerns the first listed respondent. The nature of this litigant falls into the category "federal government (including DC)", specifically "cabinet level department". Which specific federal government agency best describes this litigant? Department of Agriculture, Department of Commerce, Department of Defense (includes War Department and Navy Department)
songer_respond1_4_2 This question concerns the first listed respondent. The nature of this litigant falls into the category "sub-state government (e.g., county, local, special district)". Which category of substate government best describes this litigant? legislative, executive/administrative, bureaucracy providing services
songer_respond1_4_3 This question concerns the first listed respondent. The nature of this litigant falls into the category "sub-state government (e.g., county, local, special district)", specifically "legislative". Which specific substate government agency best describes this litigant? City/county council, School Board, board of trustees for college or junior college, Other legislative body
songer_respond1_5_2 This question concerns the first listed respondent. The nature of this litigant falls into the category "state government (includes territories & commonwealths)". Which category of state government best describes this litigant? legislative, executive/administrative, bureaucracy providing services
songer_respond1_5_3 This question concerns the first listed respondent. The nature of this litigant falls into the category "state government (includes territories & commonwealths)", specifically "legislative". Which specific state government agency best describes this litigant? Legislature or separate house as an organization, Legislative Committee or Commission, Other Legislative Unit
songer_respond1_7_2 This question concerns the first listed respondent. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the gender of this litigant?Use names to classify the party’s sex only if there is little ambiguity. not ascertained, male - indication in opinion (e.g., use of masculine pronoun), male - assumed because of name
songer_respond1_7_3 This question concerns the first listed respondent. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the race or ethnic identity of this litigant as identified in the opinion? not ascertained, caucasian - specific indication in opinion, black - specific indication in opinion
songer_respond1_7_4 This question concerns the first listed respondent. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the citizenship of this litigant as indicated in the opinion? not ascertained, US citizen, alien
songer_respond1_7_5 This question concerns the first listed respondent. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". Which of these categories best describes the income of the litigant? not ascertained, poor + wards of state, presumed poor
songer_respond1_8_2 This question concerns the first listed respondent. The nature of this litigant falls into the category "miscellaneous". Which of the following categories best describes the litigant? fiduciary, executor, or trustee, other, nature of the litigant not ascertained
songer_respond1_8_3 This question concerns the first listed respondent. The nature of this litigant falls into the category "miscellaneous", specifically "fiduciary, executor, or trustee". Which of the following specific subcategories best describes the litigant? trustee in bankruptcy - institution, trustee in bankruptcy - individual, executor or administrator of estate - institution
songer_respond2_1_2 This question concerns the second listed respondent. The nature of this litigant falls into the category "private business (including criminal enterprises)". What is the scope of this business? local, neither local nor national, national or multi-national
songer_respond2_1_3 This question concerns the second listed respondent. The nature of this litigant falls into the category "private business (including criminal enterprises)". What category of business best describes the area of activity of this litigant which is involved in this case? agriculture, mining, construction
songer_respond2_1_4 This question concerns the second listed respondent. The nature of this litigant falls into the category "private business (including criminal enterprises)", specifically "agriculture". What subcategory of business best describes this litigant? single family farm, commercial farm, agri-business, farm - other
songer_respond2_2_2 This question concerns the second listed respondent. The nature of this litigant falls into the category "private organization or association". What category of private associations best describes this litigant? business, trade, professional, or union (BTPU), other
songer_respond2_2_3 This question concerns the second listed respondent. The nature of this litigant falls into the category "private organization or association", specifically "business, trade, professional, or union (BTPU)". What subcategory of private association best describes this litigant? Business or trade association, utilities co-ops, Professional association - other than law or medicine
songer_respond2_3_2 This question concerns the second listed respondent. The nature of this litigant falls into the category "federal government (including DC)". Which category of federal government agencies and activities best describes this litigant? cabinet level department, courts or legislative, agency whose first word is "federal"
songer_respond2_3_3 This question concerns the second listed respondent. The nature of this litigant falls into the category "federal government (including DC)", specifically "cabinet level department". Which specific federal government agency best describes this litigant? Department of Agriculture, Department of Commerce, Department of Defense (includes War Department and Navy Department)
songer_respond2_4_2 This question concerns the second listed respondent. The nature of this litigant falls into the category "sub-state government (e.g., county, local, special district)". Which category of substate government best describes this litigant? legislative, executive/administrative, bureaucracy providing services
songer_respond2_4_3 This question concerns the second listed respondent. The nature of this litigant falls into the category "sub-state government (e.g., county, local, special district)", specifically "legislative". Which specific substate government agency best describes this litigant? City/county council, School Board, board of trustees for college or junior college, Other legislative body
songer_respond2_5_2 This question concerns the second listed respondent. The nature of this litigant falls into the category "state government (includes territories & commonwealths)". Which category of state government best describes this litigant? legislative, executive/administrative, bureaucracy providing services
songer_respond2_5_3 This question concerns the second listed respondent. The nature of this litigant falls into the category "state government (includes territories & commonwealths)", specifically "legislative". Which specific state government agency best describes this litigant? Legislature or separate house as an organization, Legislative Committee or Commission, Other Legislative Unit
songer_respond2_7_2 This question concerns the second listed respondent. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the gender of this litigant?Use names to classify the party’s sex only if there is little ambiguity. not ascertained, male - indication in opinion (e.g., use of masculine pronoun), male - assumed because of name
songer_respond2_7_3 This question concerns the second listed respondent. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the race or ethnic identity of this litigant as identified in the opinion? not ascertained, caucasian - specific indication in opinion, black - specific indication in opinion
songer_respond2_7_4 This question concerns the second listed respondent. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". What is the citizenship of this litigant as indicated in the opinion? not ascertained, US citizen, alien
songer_respond2_7_5 This question concerns the second listed respondent. The nature of this litigant falls into the category "natural person (excludes persons named in their official capacity or who appear because of a role in a private organization)". Which of these categories best describes the income of the litigant? not ascertained, poor + wards of state, presumed poor
songer_respond2_8_2 This question concerns the second listed respondent. The nature of this litigant falls into the category "miscellaneous". Which of the following categories best describes the litigant? fiduciary, executor, or trustee, other, nature of the litigant not ascertained
songer_respond2_8_3 This question concerns the second listed respondent. The nature of this litigant falls into the category "miscellaneous", specifically "fiduciary, executor, or trustee". Which of the following specific subcategories best describes the litigant? trustee in bankruptcy - institution, trustee in bankruptcy - individual, executor or administrator of estate - institution
songer_rtcouns Did the court rule that the defendant’s right to counsel was violated (for some reason other than inadequate counsel)? No, Yes, Yes, but error was harmless
songer_search Did the court below improperly rule for the prosecution on an issue related to an alleged illegal search and seizure? No, Yes, Yes, but error was harmless
songer_sentence Did the court conclude that some penalty, excluding the death penalty, was improperly imposed? No, Yes, Yes, but error was harmless
songer_source What forum heard this case immediately before the case came to the court of appeals? Federal district court (single judge), 3 judge district court, State court
songer_st_v_st Did the court rule in favor of the appellant on the issue of a conflict of laws ( which laws or rules apply ) other than federal v state or foreign v domestic (e.g., one state vs second state)? No, Yes, Mixed answer
songer_standing Did the court determine that the parties had standing? No, Yes, Mixed answer
songer_state In what state or territory was the case first heard? not, Alabama, Alaska
songer_stateclaim Did the court dismiss the case because of the failure of the plaintiff to state a claim upon which relief could be granted? No, Yes, Mixed answer
songer_stpolicy Did the interpretation of state or local law, executive order, administrative regulation, doctrine, or rule of procedure by the court favor the appellant? No, Yes, Mixed answer
songer_subevid Did the court’s interpretation of the substantial evidence rule support the government? For example, "such evidence as a reasonable mind might accept as adequate to support a conclusion" or "more than a mere scintilla". This issue is present only when the court indicates that it is using this doctrine, rather than when the court is merely discussing the evidence to determine whether the evidence supports the position of the appellant or respondent. No, Yes, Mixed answer
songer_suffic Did the court rule that there was insufficient evidence for conviction? No, Yes, Yes, but error was harmless
songer_summary Did the court’s ruling on the appropriateness of summary judgment or the denial of summary judgment favor the appellant? No, Yes, Mixed answer
songer_timely Did the court conclude that it could not reach the merits of the case because the litigants had not complied with some rule relating to timeliness, a filing fee, or because a statute of limitations had expired? No, Yes, Mixed answer
songer_treat What is the disposition by the court of appeals of the decision of the court or agency below? stay, petition, or motion granted, affirmed; or affirmed and petition denied, reversed (include reversed & vacated)
songer_trialpro Did the court’s ruling on procedure at trial favor the appellant? This includes jury instructions and motions for directed verdicts made during trial. No, Yes, Mixed answer
songer_two_issues Are there two issues in the case? no, yes
songer_typeiss What is the general category of issues discussed in the opinion of the court? criminal and prisoner petitions, civil - government, diversity of citizenship
songer_usc1 What is the most frequently cited title of the U.S. Code in the headnotes to this case? Answer with a number. N/A
songer_usc1sect What is the number of the section from the title of the most frequently cited title of the U.S. Code in the headnotes to this case, that is, title usc1? Answer with a number. N/A
songer_usc2 The most frequently cited title of the U.S. Code in the headnotes to this case is usc1. What is the second most frequently cited title of this U.S. Code in the headnotes to this case? Answer with a number. N/A
songer_usc2sect What is the number of the section from the title of the second most frequently cited title of the U.S. Code in the headnotes to this case, that is, title usc2? Answer with a number. N/A
songer_weightev Did the factual interpretation by the court or its conclusions (e.g., regarding the weight of evidence or the sufficiency of evidence) favor the appellant? No, Yes, Mixed answer
songer_whlaws Did the court’s discussion of which state’s laws should control their ruling in the case support the position taken by the appellant? No, Yes, Mixed answer